# Mamba-in-Mamba: Centralized Mamba-Cross-Scan in Tokenized Mamba Model for Hyperspectral Image Classification

### v3

@@ Abstract
Hyperspectral image (HSI) classification is pivotal in the remote sensing (RS) field, particularly with the advancement of deep learning techniques. Sequential models, adapted from the natural language processing (NLP) field—such as Recurrent Neural Networks (RNNs) and Transformers—have been tailored to this task, offering unique viewpoints. However, they face several problems: 1) RNNs struggle with aggregating central features and are sensitive to interfering pixels; 2) Transformers require extensive computational resources and often underperform with limited HSI training samples. To address the issues associated with RNNs and Transformers, recent advances have introduced State Space Models (SSM) and Mamba, known for their lightweight and parallel scanning capabilities within linear sequence processing, providing a balance between the two. However, the application of existing Mamba-based models to HSI classification has not been fully explored, and the suitability of these models for this task remains an open question. Furthermore, the current methods for converting images into sequence data in Mamba-based models are simplistic and inefficient. In response, this study introduces the innovative Mamba-in-Mamba (MiM) architecture for HSI classification, marking the pioneering deployment of Mamba in this field. The MiM model includes: 1) A novel centralized Mamba-Cross-Scan (MCS) mechanism for transforming images into sequence data; 2) A Tokenized Mamba (T-Mamba) encoder that incorporates a Gaussian Decay Mask (GDM), a Semantic Token Learner (STL), and a Semantic Token Fuser (STF) for enhanced feature generation and concentration; and 3) A Weighted MCS Fusion (WMF) module coupled with a Multi-Scale Loss Design to improve decoding efficiency. Experimental results from four public HSI datasets with fixed and disjoint training-testing samples demonstrate that our method outperforms existing baselines and state-of-the-art approaches, highlighting its efficacy and potential in HSI applications.
Index Terms:
@@
Hyperspectral image classification, deep learning, multiscanning, state space model, Mamba.
I Introduction
@@
Hyperspectral imaging (HSI) is a sophisticated remote sensing (RS) technique that captures images across numerous dense and continuous spectral bands. This method spans the electromagnetic spectrum from visible to infrared wavelengths, often utilizing satellites or unmanned aerial vehicles (UAVs) [1]. Unlike traditional imaging methods, HSI provides detailed spectral information for each pixel, enabling a more precise characterization of materials [2] and objects on the Earth’s surface [3]. The comprehensive data obtained from HSI makes it invaluable in various applications including agriculture [4], environmental monitoring [5], geology [6], urban planning [7], and defense and security [8]. Given the broad utility of this technology, accurate HSI classification, which assigns a label to each pixel based on its spectral signature, is critical. This classification task is not only essential for enhancing the effectiveness of these applications but has also become a focus of research within the RS community [9].
@@
In recent years, the deployment of deep learning-based models for HSI classification has seen extensive growth [10]. These models typically employ a ‘patch-wise’ learning framework, where a local patch around each pixel is processed to assign semantic labels [11, 12]. Notably, sequential models like Recurrent Neural Networks (RNNs) [13, 14, 15, 16, 17] and Transformers [18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28], originally developed for natural language processing (NLP), have been adapted to HSI classification. They have demonstrated impressive results by utilizing many-to-one or many-to-many schemes to capture contextual features within pixels or partitioned windows of one HSI patch, resulting in a comprehensive feature representation. Despite their success, both types of models face shortcomings that warrant further investigation to enhance their practicability and efficiency.
Second, to address the inflexibility of RNNs in HSI classification, Transformer-based methods have been introduced, utilizing parallel self-attention to enhance performance [35]. While these methods overcome some of RNNs’ limitations through parallel processing and self-attention mechanisms, they hold their own set of shortcomings: 1) The self-attention mechanism scales quadratically with the size of HSI patches, leading to significant computational overhead. This becomes especially problematic when larger patches are used to capture more spatial or contextual information around the center pixel, making the approach resource-intensive and less suited for hardware-constrained environments; 2) Transformers often have a large number of parameters, increasing the risk of overfitting, especially with the limited training data typically available in HSI classification; 3) Unlike RNNs or convolutional neural networks (CNNs) [36, 37, 38], which inherently capture recurrent features or local spatial features, Transformers do not have this inductive bias. While positional embeddings are used to introduce some spatial context, their effectiveness is limited without substantial training samples, potentially reducing the model’s ability to accurately capture spatial contexts in HSIs. Although some research [39, 40, 41, 42] has explored hybrid approaches that integrate RNNs or CNNs with Transformers to leverage complementary strengths. However, these methods may still increase the computational burden and complexity, potentially leading to suboptimal performance.
In response to three scenarios discussed, two critical needs emerge: 1) Alternative Base Model: There is a need for an alternative model that enhances parallel computing or incorporates attention mechanisms, offering greater flexibility than RNNs. This model should also demand fewer resources and exhibit less complexity, proving lighter and more efficient than Transformers, and 2) Enhanced Scanning Method: A refined scanning method is required that supports continuous, multi-directional processing. This method should provide image-aware cognition without being redundant, effectively capturing the spatial correlations and dynamics inherent in image data.
Recently, structured state space sequence (S4) models [46] and the Mamba architecture [47, 48, 49, 50] are brought into the spotlight. These models are adept at handling long-sequence data and can scale linearly with sequence length. The Mamba architecture, in particular, builds on S4 by adding a selective-scan mechanism that adaptively attentions on relevant inputs, showing superior computational efficiency compared to Transformers. Although the Mamba-based methods has shown promising results in some visual tasks [51, 52, 54, 53, 55, 56] (e.g. Vision Mamba (ViM), Visual Mamba (VMamba), etc), its potential for HSI classification has yet to be fully explored. Therefore, there is a critical need to develop a suitable vision-based Mamba model specifically designed for HSI classification.
@@
    We explore the potential applications of the improved ViM architecture for HSI classification task thereby achieving the better performances and efficiency.
@@
    We propose centralized MCS to comprehensively and continuously scan the HSI patch into several sequences diversely, which can fully and lightly understand the HSI patch within the highlights on the center pixel.
@@
    The proposal shows very competitive and even SOTA performance on four datasets with fixed and disjoint training-testing samples, demonstrating the feasibility.
The rest of the paper is organized as follows: Section II presents the preliminary knowledge of Mamba, Section III details the methodology, Section IV discusses the experimental setup and results, and Section V concludes the study.
II Preliminary
@@
II-A Structured State Space Sequence Model (S4)
@@
II-B Selective Scan-Based S4 Model (S6)
@@
Algorithm 1 Pseudo procedure of the proposed method
2:  Output: Predicted labels of the testing samples.
4:  Procedure:
9:  Apply four types of MCS on $\textbf{X}^{(i,\,j)}$, getting sequences, $\mathbf{S}_{1}^{(i,\,j)}$, $\mathbf{S}_{2}^{(i,\,j)}$, $\mathbf{S}_{3}^{(i,\,j)}$, and $\mathbf{S}_{4}^{(i,\,j)}$ (Eqs. 14a-19a).
10:  MiM model:
11:  for $m=1$ to $4$ in $\mathbf{S}_{m}^{(i,\,j)}$ do
13:     T-Mamba encoder:
14:     for Both forward pass $\mathbf{s}_{f}$ and backward pass $\mathbf{s}_{b}$ do
21:        Activate $\mathbf{o}_{m}^{(i,\,j)}$ using the Tanh function.
22:     end for
23:  end for
II-C Mamba-Based Works
@@
Despite these advancements, to our knowledge, the application of ViM with multi-directional scanning for HSI classification has not yet been fully developed or explored. Identifying an optimized ViM architecture for HSI remains an open question. Thus, we aim to explore this gap and propose a tailored HSI-driven ViM architecture.
Figure 3: The overall structure. It begins with a cropped HSI patch after PCA and then goes through a depth-wise convolution and a point-wise convolution for local feature generation and dynamic positional embedding. It is then divided into four types of sequences by MCS, as shown in Fig. 5. Each scan type is independently processed by a specially designed T-Mamba encoder without sharing parameters, as detailed in Fig. 6, to extract contextual features. The outputs from the four T-Mamba encoders are combined through a WMF module with dynamic weighted scan fusion. Concurrently, fused feature at current scale is recorded for final multi-scale decoding. This process is repeated, with the feature sizes being progressively down-sampled until they reach a spatial size of $1$. Finally, the accumulated multi-scale features from each iteration are compiled to feed into a decoder for model training.
III Methodology
This section firstly presents (A) methodological overview briefly. Then, it presents the detailed contents: (B) Centralized MCS, (C) T-Mamba Encoder, (D) WMF Module, (E) Multi-Scale Loss Design. The overall structure and the pseudo algorithm is shown in Fig. 3 and Algorithm 1, respectively.
III-A Methodological Overview
Our method is composed of several parts:
  1. 1.
@@ III-A Methodological Overview
Detailed explanations for each component are provided in the subsequent subsections.
III-B Centralized Mamba-Cross-Scan (MCS)
@@
Before feeding the $\mathbf{X}^{(i,\,j)}$ into MCS, we first deploy the depth-wise convolution (DWconv) [^3] https://github.com/rosinality/depthwise-conv-pytorch and point-wise convolution (PWconv) [^4] https://github.com/anshulpaigwar/Pointwise-Convolutional-Neural-Network on $\mathbf{X}^{(i,\,j)}$ for local feature aggregation and dynamic positional embedding.
$\mathbf{X}^{(i,\,j)}\leftarrow PWconv\Big{(}DWconv\Big{(}\mathbf{X}^{(i,\,j)}\Big{)}\Big{)}$ (10)
III-B1 MCS Type-1
@@
$\begin{array}[]{ccccccccc}&\overbrace{{\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\mathbf{\mathbf{x}}^{(i-2,j-2)}}}^{Start}&\rightarrow&\mathbf{\mathbf{x}}^{(i-2,j-1)}&\rightarrow&\cdots&\rightarrow\mathbf{\mathbf{x}}^{(i-2,j+2)}\\ &&&&&&\downarrow\\ &\mathbf{\mathbf{x}}^{(i-1,j-2)}&\leftarrow&\mathbf{\mathbf{x}}^{(i-1,j-1)}&\leftarrow&\cdots&\leftarrow\mathbf{\mathbf{x}}^{(i-1,j+2)}\\ &\downarrow&&&&&\\ &\mathbf{\mathbf{x}}^{(i,j-2)}&\rightarrow&\mathbf{\mathbf{x}}^{(i,j-1)}&\rightarrow&\cdots&\rightarrow\mathbf{\mathbf{x}}^{(i,j+2)}\\ &&&&&&\downarrow\\ &\mathbf{\mathbf{x}}^{(i+1,j-2)}&\leftarrow&\mathbf{\mathbf{x}}^{(i+1,j-1)}&\leftarrow&\cdots&\leftarrow\mathbf{\mathbf{x}}^{(i+1,j+2)}\\ &\downarrow&&&&&\\ &\mathbf{\mathbf{x}}^{(i+2,j-2)}&\rightarrow&\mathbf{\mathbf{x}}^{(i+2,j-1)}&\rightarrow&\cdots&\rightarrow\underbrace{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{\mathbf{x}}^{(i+2,j+2)}}}_{End}\\ \end{array}$ (11)
Therefore, the generated complete sequence $\mathbf{S}_{1}^{(i,\,j)}\in\mathbb{R}^{p^{2}\times C}$ can be represented as:
$\begin{split}\mathbf{S}^{(i,j)}_{1}=&[{\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\mathbf{x}^{(i-2,j-2)}},\mathbf{x}^{(i-2,j-1)},\cdots,\mathbf{x}^{(i-2,j+2)},\cdots\\ &\cdots,\mathbf{x}^{(i+2,j-2)},\mathbf{x}^{(i+2,j-1)},\cdots,{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{x}^{(i+2,j+2)}}]^{\top},\end{split}$ (12)
@@
$\mathbf{S}^{(i,j)}_{1,\,f}$ $=\Big{[}\underbrace{{\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\mathbf{x}^{(i-2,j-2)}},\mathbf{x}^{(i-1,j-2)},\mathbf{x}^{(i,j-2)},\cdots,{\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}\mathbf{x}^{(i,j)}}}_{{length}=(p^{2}+1)/2}\Big{]}^{\top},$ (14a)
$\mathbf{S}^{(i,j)}_{1,\,b}$ $=\Big{[}\underbrace{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{x}^{(i+2,j+2)}},\mathbf{x}^{(i+1,j+2)},\mathbf{x}^{(i,j+2)},\cdots,{\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}\mathbf{x}^{(i,j)}}}_{{length}=(p^{2}+1)/2}\Big{]}^{\top},$ (14b)
where two sub-sequences termed as forward-direction and backward-direction, both include the center step $\mathbf{x}^{(i,j)}$ of the complete sequence and assign it at the last step in the sub-sequence. In this case, when patch size $p$ equals 5, the length of sub-sequence is 13, as shown in Fig. 5(a)

### v4

@@ Abstract
Hyperspectral image (HSI) classification is pivotal in the remote sensing (RS) field, particularly with the advancement of deep learning techniques. Sequential models, such as Recurrent Neural Networks (RNNs) and Transformers, have been tailored to this task, offering unique viewpoints. However, they face several challenges: 1) RNNs struggle with aggregating central features and are sensitive to interfering pixels; 2) Transformers require extensive computational resources and often underperform with limited HSI training samples. To address these issues, recent advances have introduced State Space Models (SSM) and Mamba, known for their lightweight and parallel scanning capabilities within linear sequence processing, providing a balance between RNNs and Transformers. However, deploying Mamba as a backbone for HSI classification has not been fully explored. Although there are improved Mamba models for visual tasks, such as Vision Mamba (ViM) and Visual Mamba (VMamba), directly applying them to HSI classification encounters problems. For example, these models do not effectively handle land-cover semantic tokens with multi-scale perception for feature aggregation when implementing patch-based HSI classifiers for central pixel prediction. Hence, the suitability of these models for this task remains an open question. In response, this study introduces the innovative Mamba-in-Mamba (MiM) architecture for HSI classification, marking the pioneering deployment of Mamba in this field. The MiM model includes: 1) A novel centralized Mamba-Cross-Scan (MCS) mechanism for transforming images into efficient-pair sequence data; 2) A Tokenized Mamba (T-Mamba) encoder that incorporates a Gaussian Decay Mask (GDM), a Semantic Token Learner (STL), and a Semantic Token Fuser (STF) for enhanced feature generation and concentration; and 3) A Weighted MCS Fusion (WMF) module coupled with a Multi-Scale Loss Design to improve model training efficiency. Experimental results from four public HSI datasets with fixed and disjoint training-testing samples demonstrate that our method outperforms existing baselines and is competitive with state-of-the-art approaches, highlighting its efficacy and potential in HSI applications.
keywords:
@@
Hyperspectral image classification, deep learning, state space model, Mamba.
[…]
{graphicalabstract}
{highlights}
We explore the potential design of Mamba architecture, Mamba-in-Mamba model, for HSI classification task thereby achieving the better competitive performances and satisfying efficiency.
To fully explore the ability of Mamba in visual task, we propose a novel Tokenized Mamba encoder with Gaussian Decay Mask, Semantic Token Learner, and Semantic Token Fuser to improve its suitability, aligning with multi-scale features for the HSI classification.
We propose a centralized Mamba-Cross-Scan to comprehensively and continuously scan the image patch into several pair-sequences diversely. It can fully and lightly understand the HSI patch within the concentration on the center pixel for patch-wise HSI classifiers.
This study shows very competitive and even SOTA performance on four datasets, Indian Pines, Pavia University, Houston 2013, and WHU-Hi-HongHu, with fixed and disjoint training-testing samples, demonstrating the feasibility and efficiency.
1 Introduction
@@
Hyperspectral imaging (HSI) is a sophisticated remote sensing (RS) technique that captures images across numerous dense and continuous spectral bands. This method spans the electromagnetic spectrum from visible to infrared wavelengths, often utilizing satellites or unmanned aerial vehicles (UAVs) [1]. Unlike traditional imaging methods, HSI provides detailed spectral information for each pixel, enabling a more precise characterization of materials [2] and objects on the Earth’s surface [3]. The comprehensive data obtained from HSI makes it invaluable in various applications including agriculture [4], environmental monitoring [5], geology [6], urban planning [7], and defense and security [8]. Given the broad utility of this technology, accurate HSI classification, which assigns a land-cover label to each pixel, is critical. This classification task is not only essential for enhancing the effectiveness of these applications but has also become a focus of research within the RS community [9].
In recent years, the deployment of deep learning-based models for HSI classification has seen extensive growth [10]. These models typically employ a ‘patch-wise’ learning framework, where a local patch around each pixel is processed to assign semantic labels [11, 12]. Notably, sequential models like Recurrent Neural Networks (RNNs) [13, 14, 15, 16, 17] and Transformers [18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33], originally developed for natural language processing (NLP), have been adapted to HSI classification. They have demonstrated impressive results by utilizing many-to-one or many-to-many schemes to capture contextual features within pixels or partitioned windows of one HSI patch, resulting in a comprehensive feature representation. Despite their success, both types of models face shortcomings that warrant further investigation to enhance their practicability and efficiency.
Second, to address the inflexibility of RNNs, Transformer-based methods have been introduced, utilizing parallel self-attention to enhance performance [34]. While these methods overcome some of RNNs’ limitations through parallel processing and self-attention mechanisms, they hold their own set of shortcomings: 1) The self-attention mechanism scales quadratically with the size of HSI patches, leading to significant computational overhead. This becomes especially problematic when larger patches are used to capture more spatial or contextual information around the center pixel, making the approach resource-intensive and less suited for hardware-constrained environments; 2) Transformers often have a large number of parameters, increasing the risk of overfitting, especially with the limited training data typically available in HSI classification; 3) Unlike RNNs or convolutional neural networks (CNNs) [35, 36, 37], which inherently capture recurrent features or local spatial features, Transformers do not have this inductive bias. While positional embeddings are used to introduce some spatial context, their effectiveness is limited without substantial training samples, potentially reducing the model’s ability to accurately capture spatial contexts in HSIs. Although some research [38, 39, 40, 41] has explored hybrid approaches that integrate RNNs or CNNs with Transformers to leverage complementary strengths. However, these methods may still increase the computational burden and complexity, potentially leading to suboptimal performance.
In response to three scenarios discussed above, two critical needs emerge: 1) Alternative Base Model: There is a need for an alternative model that enhances parallel computing or incorporates attention mechanisms, offering greater flexibility than RNNs. This model should also demand fewer resources and exhibit less complexity, proving lighter and more efficient than Transformers, and 2) Enhanced Scanning Method: A refined scanning method is required that supports continuous and multi-directional processing. This method should provide image-aware cognition without being redundant, effectively capturing the spatial correlations and dynamics inherent in image data.
Recently, structured state space sequence (S4) models [45] and the Mamba architecture [46, 47, 48, 49] are brought into the spotlight. These models are adept at handling long-sequence data and can scale linearly with sequence length. The Mamba architecture, in particular, builds on S4 by adding a selective-scan mechanism that adaptively attentions on relevant inputs, showing superior computational efficiency compared to Transformers. Although the Mamba-based methods has shown promising results in some visual tasks [50, 51, 53, 52, 54, 55] (e.g. Vision Mamba (ViM), Visual Mamba (VMamba), etc), its potential for HSI classification has yet to be fully explored. Directly applying these advanced Mamba-based models on HSI classification task may encounter unsuitability and problems. Therefore, there is a critical need to develop a suitable vision-based Mamba model specifically designed for HSI classification.
@@
    We explore the potential applications of the improved ViM architecture, multi-scale MiM model, for HSI classification task thereby achieving the better performances and efficiency.
@@
    We propose centralized MCS to comprehensively and continuously scan the HSI patch into several pair-sequences diversely, which can fully and lightly understand the HSI patch within the highlights on the center pixel for patch-wise HSI classifiers.
@@
    The proposal shows very competitive and even SOTA performance on four datasets with fixed and disjoint training-testing samples, demonstrating the feasibility and efficiency.
The rest of the paper is organized as follows: Section 2 presents the preliminary knowledge of Mamba, Section 3 details the methodology, Section 4 discusses the experimental setup and results, and Section 5 concludes the study.
2 Preliminary
@@
2.1 Structured State Space Sequence Model (S4)
@@
2.2 Selective Scan-Based S4 Model (S6)
@@
2.3 Mamba-Based Works
@@
Despite these advancements, to our knowledge, the application of ViM with multi-directional scanning for HSI classification has not yet been fully developed or explored. Meanwhile, directly using ViM can not well handle the refined semantic token features and miss the multi-scale features existed in the HSI patch. Hence, developing an optimized ViM architecture for HSI remains an open question. Thus, we aim to explore this gap and propose a tailored HSI-driven ViM architecture.
Figure 3: The overall structure. It begins with a cropped HSI patch after PCA and then goes through a depth-wise convolution and a point-wise convolution for local feature generation and dynamic positional embedding. It is the reshaped by snake flattening before MiM model. In the MiM model, flattened image is first proposed by MCS, as shown in Fig. 5, to obtain four types of pair-sequences. Each pair of sequences processed by a specially designed T-Mamba encoder with sharing parameters, as detailed in Fig. 6, to extract contextual features. The outputs from the four T-Mamba encoders are combined through a WMF module with dynamic weighted scan fusion. Concurrently, fused feature at current scale is recorded for final multi-scale decoding. This process is repeated, with the feature sizes being progressively down-sampled until they reach a spatial size of $1$. Finally, the accumulated multi-scale features from each MiM iteration are compiled to feed into a decoder for model training and decoding.
3 Methodology
This section firstly presents (A) methodological overview briefly. Then, it presents the detailed contents: (B) Centralized MCS, (C) T-Mamba Encoder, (D) WMF Module, (E) Multi-Scale Loss Design. The overall structure and the pseudo algorithm is shown in Fig. 3 and Algorithm 1, respectively. Due to the paper formatting, the summary of algorithm is placed at end of the article.
3.1 Methodological Overview
Our method is composed of several following steps:
  1. 1)
@@ 3.1 Methodological Overview
Detailed explanations for each component are provided in the subsequent subsections later.
3.2 Centralized Mamba-Cross-Scan (MCS)
@@
Before feeding the $\mathbf{X}^{(i,\,j)}$ into MCS, we first deploy the principal component analysis (PCA), depth-wise convolution (DWconv) [^3] https://github.com/rosinality/depthwise-conv-pytorch and point-wise convolution (PWconv) [^4] https://github.com/anshulpaigwar/Pointwise-Convolutional-Neural-Network on $\mathbf{X}^{(i,\,j)}$ for dimension reduction, local feature aggregation and dynamic positional embedding, as:
$\mathbf{X}^{(i,\,j)}\leftarrow PWconv\Big{(}DWconv\Big{(}PCA\Big{(}\mathbf{X}^{(i,\,j)}\Big{)}\Big{)}\Big{)}$ (10)
where the PCA preserves the number of $\mathcal{C}_{1}$ components. The PWconv and DWconv outputs number of $\mathcal{C}_{2}$ feature dimensions, where $\mathcal{C}_{2}>\mathcal{C}_{1}$. Hence, the resultant $\mathbf{X}^{(i,\,j)}\in\mathbb{R}^{p\times p\times\mathcal{C}_{2}}$.
3.2.1 MCS Type-1
@@
$\begin{array}[]{ccccccccc}&\overbrace{{\color[rgb]{0,1,0}\mathbf{\mathbf{x}}^{(i-2,j-2)}}}^{Start}&\rightarrow&\mathbf{\mathbf{x}}^{(i-2,j-1)}&\rightarrow&\cdots&\rightarrow\mathbf{\mathbf{x}}^{(i-2,j+2)}\\ &&&&&&\downarrow\\ &\mathbf{\mathbf{x}}^{(i-1,j-2)}&\leftarrow&\mathbf{\mathbf{x}}^{(i-1,j-1)}&\leftarrow&\cdots&\leftarrow\mathbf{\mathbf{x}}^{(i-1,j+2)}\\ &\downarrow&&&&&\\ &\mathbf{\mathbf{x}}^{(i,j-2)}&\rightarrow&\mathbf{\mathbf{x}}^{(i,j-1)}&\rightarrow&\cdots&\rightarrow\mathbf{\mathbf{x}}^{(i,j+2)}\\ &&&&&&\downarrow\\ &\mathbf{\mathbf{x}}^{(i+1,j-2)}&\leftarrow&\mathbf{\mathbf{x}}^{(i+1,j-1)}&\leftarrow&\cdots&\leftarrow\mathbf{\mathbf{x}}^{(i+1,j+2)}\\ &\downarrow&&&&&\\ &\mathbf{\mathbf{x}}^{(i+2,j-2)}&\rightarrow&\mathbf{\mathbf{x}}^{(i+2,j-1)}&\rightarrow&\cdots&\rightarrow\underbrace{{\color[rgb]{0,0,1}\mathbf{\mathbf{x}}^{(i+2,j+2)}}}_{End}\\ \end{array}$ (11)
Therefore, the generated complete sequence $\mathbf{S}_{1}^{(i,\,j)}\in\mathbb{R}^{p^{2}\times\mathcal{C}_{2}}$ can be represented as:
$\begin{split}\mathbf{S}^{(i,j)}_{1}=&[{\color[rgb]{0,1,0}\mathbf{x}^{(i-2,j-2)}},\mathbf{x}^{(i-2,j-1)},\cdots,\mathbf{x}^{(i-2,j+2)},\cdots\\ &\cdots,\mathbf{x}^{(i+2,j-2)},\mathbf{x}^{(i+2,j-1)},\cdots,{\color[rgb]{0,0,1}\mathbf{x}^{(i+2,j+2)}}]^{\top},\end{split}$ (12)
@@
$\mathbf{S}^{(i,j)}_{1,\,f}$ $=\Big{[}\underbrace{{\color[rgb]{0,1,0}\mathbf{x}^{(i-2,j-2)}},\mathbf{x}^{(i-1,j-2)},\mathbf{x}^{(i,j-2)},\cdots,{\color[rgb]{1,.5,0}\mathbf{x}^{(i,j)}}}_{{length}=(p^{2}+1)/2}\Big{]}^{\top},$ (14a)
$\mathbf{S}^{(i,j)}_{1,\,b}$ $=\Big{[}\underbrace{{\color[rgb]{0,0,1}\mathbf{x}^{(i+2,j+2)}},\mathbf{x}^{(i+1,j+2)},\mathbf{x}^{(i,j+2)},\cdots,{\color[rgb]{1,.5,0}\mathbf{x}^{(i,j)}}}_{{length}=(p^{2}+1)/2}\Big{]}^{\top},$ (14b)
where two sub-sequences termed as forward-direction and backward-direction, both include the center step $\mathbf{x}^{(i,j)}$ of the complete sequence and assign it at the last step in the sub-sequence. In this case, when patch size $p$ equals 5, the length of sub-sequence is 13, as shown in Fig. 5(a).